OpenStudy (anonymous):

Solve the system. y = 4x + 12 and y = -2x + 6 (0, 3) (-1, 8) (12, 6) no solution infinitely many solutions ok so what i know is im suppose to write it like this right? 4x+12=-2x+6 what next?

OpenStudy (saifoo.khan):

Yes, you're right. Solve for x.

OpenStudy (anonymous):

4x+12/4=-2x+6/4 would it be like that next? im not sure

OpenStudy (saifoo.khan):

No. That will make this complicated. bring variable on one side of the equation.

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

6x=-6?

OpenStudy (saifoo.khan):

Yes. Now we're going to do next?

OpenStudy (anonymous):

6x/6=-6/6 x=-1?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

sry.. no solution

OpenStudy (anonymous):

hu?

OpenStudy (saifoo.khan):

Correct! Now insert this value in equation 1 or 2 to get the value of y.

OpenStudy (anonymous):

y=4(-1)+12 y=-4+12 y=8

OpenStudy (anonymous):

so -1,8 ?

OpenStudy (precal):